OpenStudy (anonymous):

what is the anti-derivative of this function? I screwed it up somewhere and I am not sure of my answer. 1.82(.89^X)

myininaya (myininaya):

\[\int\limits_{}^{}a ^x dx=\frac{a^x}{\ln(a)} +C \text{ for a>0}\]

OpenStudy (anonymous):

if the data started at 1982 and ended at 2000, does a=0 and b=18 if you wanted to know the definite integral between that interval?